what is the measure of the smaller angle of two supplementary angle if twice the larger is ten less than three times the small angle.
Answer by stanbon(75887)

what is the measure of the smaller angle of two supplementary angle if twice the larger is ten less than three times the small angle.
-----
Let the smaller angle be "x".
Then its supplement is "180-x".
---
Equation:
2(180-x) = 3x-10
---
360-2x = 3x-10
5x = 370
x = 74 degrees (the smaller angle)
---
180-x = 106 degrees (the larger angle)
=========================================
Cheers,
Stan H.
==========